Jack holds a portfolio with the following securities: Security Investment Return Stock A 101,128 11.2% Stock B 364,251 4.1% Stock C 602,302 6.7% Calculate the expected return of portfolio.

Round the answers to two decimal places in percentage form.
Answer #1

Total value = 101,128 + 364,251 + 602,302 = 1,067,681

Expected return = (101,128 / 1,067,681)*0.112 + (364,251 / 1,067,681)*0.041 + (602,302 / 1,067,681)*0.067

Expected return = 0.010608 + 0.013988 + 0.037796

Expected return = 0.0624 or 6.24%